# Juding the Judges: A Systematic Investigation of Position Bias in Pairwise Comparative Assessments by LLMs

### v6

@@ C.1 Mutual Agreement & Familial Property
We compute the LLM judges’ mutual agreement on the instances to explore how “alike” or consistent they are across a set of evaluations. We denote two judges agree on an instance if their judgment choices are identical. Then the mutual agreement between two judges on a benchmark is defined as the proportion of their agreed instances. Fig. 5(c)(e) displays the mutual agreement heatmap for all judges on MTBench and DevBench, respectively. For MTBench that utilizes the 3-option mode, we also consider the “without tie” agreement since two judges are less disagreed when one chooses {C} while the other prefers a certain solution, compared to the case when they prefer different solutions. The “without tie” agreement heatmap of the twelve judges on MTBench is explored in Fig. 5(d).
@@
Gemini models exhibit rather low mutual agreement and “familial property” is minimal, but the most capable Gemini-1.5-pro aligns more closely with other capable models like the GPT-4 series and Claude-3-Opus.
@@ C.2 Disagreement & Benchmark Design Insight
Since the mutual agreement between LLM judges is not perfect and usually a considerable proportion of instances are difficult for them to reach a consensus, disagreement analysis becomes crucial and insightful. Therefore, we define the disagreement of an evaluation instance to be the number of judgments different from the majority. By this definition, an instance with all judges reaching a consensus on the better solution will have a disagreement of 0; in contrast, an instance where judgments are widely varied will result in a high disagreement. For our study where twelve judges are investigated, the maximum disagreement of an MTBench is 8, accounting for the 4{A}-4{B}-4{C} choice pattern by 3-option mode. On the other hand, for DevBench instances, the maximum possible disagreement is 6, representing the 6{A}-6{B} judgment distribution for the 2-option mode.
@@
Appendix D More Discussions on Answer Quality Gap
Intuitively, answer pairs with larger quality gaps are easier for LLM judges to evaluate, as one answer is clearly superior. Therefore, we expect LLM judges to exhibit weaker position bias in such cases. This motivates us to quantify the answer quality gap and investigate its impact on position bias. The quantification utilizes the overall win rate ($owr$) of a candidate model (i.e., answer provider) on the other to compute the quality gap. For a set of questions on a certain task and the corresponding solutions by two candidate models, an LLM judge’s evaluations would pick out the candidate winner. For the position consistent judgment instances, we can count the number of a candidate’s wins and ties, denoted as $C_{w}$ and $C_{t}$ respectively. On the other hand, following the methodology introduced by Zheng et al. (2024b), we consider the inconsistent judgment pairs as a tie for both candidates, whose counting number is denoted as $C_{I}$. This follows the intuition that it is difficult to strictly pick out a better one between two responses with similar quality, hence increasing the likelihood of position bias.
Therefore, considering both cases, we define the overall win rate ($owr$) of a candidate’s solution over the other as $owr=\frac{1}{n}[{C_{w}+\frac{1}{2}(C_{t}+C_{I})}]$, where $n$ is the total number of judgment pairs involving the two candidate solutions in both original and swapped settings. Then the quality gap ($\delta_{q}$) is calculated as $\delta_{q}=|owr-0.5|$, as 0.5 is the exact tie case for the win rate ranging from 0 to 1. We prefer the overall win rate over the consistent win rate (calculated by $\frac{C_{w}}{n_{C}}$, where $n_{C}$ is the number of position consistent judgment pairs) to quantify $\delta_{q}$ to mainly address the data sparsity issue. If the proportion of position inconsistent judgments becomes high, the available data becomes sparse and $n_{C}$ becomes small. In the extreme case, where all judgments are inconsistent, $n_{C}$ is 0 and the consistent win rate is not even computable, making the measurement inaccurate and impractical. In contrast, the overall win rate effectively captures the “comparable quality” scenario by considering “inconsistency-as-tie” and ensuring all data points contribute to the quantification of the difference between answer pairs in quality.
@@ Appendix E More Results of Position Bias and Answer Quality Gap Measurement
As shown in Fig. 4(a) and (b), considering all judges together, a larger answer quality gap generally leads to better positional consistency and fairness. In this session, we explore whether the discovery is consistent for each individual judge. Same as Section 2.2, we apply the overall win rate to reflect the answer quality gap for visualization.
@@
As shown in Fig. 6 (a), the “parabolic shape” is observed for all individual judges, indicating that the argument “a higher answer quality gap generally results in higher positional consistency” applies to all models. However, Fig. 6 (b) reveals that the positional fairness is more judge-dependent and the impact of the answer quality gap is neglectable for certain judges. For example, while Claude-3-opus and Claude-3-sonnet exhibit conspicuous “parabolic shape”, GPT-4 and GPT-3.5 present nearly linear curves. In other words, while the former models align with the general tendency that a larger answer quality gap improves positional fairness, the latter ones preserve fairness regardless of the answer quality gap. This further demonstrates the necessity to investigate positional fairness in addition to consistency when evaluating a judge model’s position bias.
@@ F.1 Bidirectional Stepwise Regression with AIC
$\log(L)=-\frac{n}{2}\log(2\pi)-\frac{n}{2}\log(\sigma^{2})-\frac{1}{2\sigma^{2}}\sum_{i=1}^{n}\hat{e}_{i}^{2},$ (5)
@@
$\text{AIC}=2k+n\log(2\pi)+n\log(\sigma^{2})+\frac{1}{\sigma^{2}}\sum_{i=1}^{n}\hat{e}_{i}^{2}.$ (6)
@@ F.2 Test results
Appendix G DevBench
Appendix H Experiment Settings
@@
H.1 Judges, Candidates, and Tasks
@@
The reference (or baseline) answer-generating models are vicuna-13b-v1.3 for MTBench and human for DevBench. They are chosen to ensure a baseline quality of responses and an expected widely spread quality gap across evaluations. The other models that are compared to the reference models, namely “Model” in our context, are listed as follows.
@@
The model names are exactly what MTBench Zheng et al. (2024b) and DevBench Li et al. (2024) used in their studies. That is why for GPTs, DevBench specifies the exact version (e.g., gpt-4-0613) while MTBench doesn’t (e.g., gpt-4). In this study, we directly use the provided answers of these models to the task questions to form answer pairs and queries for the LLM judges.
@@ H.2 Prompt Settings
  • Emphasized “should” and“shouldn’t”s.
@@
Appendix I Reproducibility
Our experiments mainly depend on API. Code will be released upon acceptance.

### v7

@@ C.1 Mutual Agreement & Familial Property
We compute the LLM judges’ mutual agreement on the instances to explore how “alike” or consistent they are across a set of evaluations. We denote two judges agree on an instance if their judgment choices are identical. Then the mutual agreement between two judges on a benchmark is defined as the proportion of their agreed instances. Fig. 4(c)(e) displays the mutual agreement heatmap for all judges on MTBench and DevBench, respectively. For MTBench that utilizes the 3-option mode, we also consider the “without tie" agreement since two judges are less disagreed when one chooses {C} while the other prefers a certain solution, compared to the case when they prefer different solutions. The “without tie” agreement heatmap of the twelve judges on MTBench is explored in Fig. 4(d).
@@
Gemini models exhibit rather low mutual agreement and “familial property" is minimal, but the most capable Gemini-1.5-pro aligns more closely with other capable models like the GPT-4 series and Claude-3-Opus.
@@ C.2 Disagreement & Benchmark Design Insight
Since the mutual agreement between LLM judges is not perfect and usually a considerable proportion of instances are difficult for them to reach a consensus, disagreement analysis becomes crucial and insightful. Therefore, we define the disagreement of an evaluation instance to be the number of judgments different from the majority. By this definition, an instance with all judges reaching a consensus on the better solution will have a disagreement of 0; in contrast, an instance where judgments are widely varied will result in a high disagreement. For our study where twelve judges are investigated, the maximum disagreement of an MTBench instance is 8, accounting for the 4{A}-4{B}-4{C} choice pattern by 3-option mode. On the other hand, for DevBench instances, the maximum possible disagreement is 6, representing the 6{A}-6{B} judgment distribution for the 2-option mode.
@@
Appendix D Open-source Model Discussion
Due to computational constraints, we conducted limited experiments on open-source models, specifically Llama 7B-13B, as representatives of smaller-scale models. However, as shown in Table 2, these models struggled to perform effectively on our evaluation tasks. A primary issue was their limited context window, which proved insufficient for benchmarks like DevBench, where questions often include extensive instructions, such as full product requirement documents. This context limitation caused frequent truncation of either the prompt or the candidate responses, leading to incomplete or erroneous outputs and significantly higher error rates.
Figure 5: Llama-13B output example: Despite providing analysis and a choice, the output format does not match the specified “[[choice]]" pattern, causing extraction failures.
Another key factor affecting performance was mismatched output formatting. MTBench requires responses in the format [[choice]] (e.g., [[A]]), but Llama-13B, as shown in Fig. 5, frequently produced outputs such as [A], which could not be parsed correctly by the regular expression used for choice extraction. Interestingly, Llama-7B outperformed Llama-13B on MTBench, likely due to its simpler, shorter responses that inadvertently aligned better with the expected format. These errors illustrate the challenges these models faced in meeting benchmark-specific requirements, highlighting nuances in their evaluation results. Considering practicality and effectiveness, we only focus on more capable commercial judge models.
@@ Appendix E More Results of Position Bias and Answer Quality Gap Measurement
E.1 MTBench
As shown in Fig. 2(c), considering all judges together, a larger answer quality gap generally leads to better position consistency and preference fairness. In this session, we explore whether the discovery is consistent for each individual judge. Same as Section 2.3, we apply the overall win rate to reflect the answer quality gap for visualization.
@@ E.1 MTBench
As shown in Fig. 6 (a), the “parabolic shape" is observed for all individual judges, indicating that the argument “a higher answer quality gap generally results in higher position consistency applies to all models. However, Fig. 6 (b) reveals that preference fairness is more judge-dependent and the impact of the answer quality gap is neglectable for certain judges. For example, while Claude-3-opus and Claude-3-sonnet exhibit conspicuous “parabolic shape", GPT-4 and GPT-3.5 present nearly linear curves. In other words, while the former models align with the general tendency that a larger answer quality gap improves preference fairness, the latter ones preserve fairness regardless of the answer quality gap. This further demonstrates the necessity to investigate preference fairness in addition to consistency when evaluating a judge model’s position bias.
E.2 DevBench
@@ F.1 Bidirectional Stepwise Regression with AIC
$\log(L)=-\frac{n}{2}\log(2\pi)-\frac{n}{2}\log(\sigma^{2})-\frac{1}{2\sigma^{2}}\sum_{i=1}^{n}\hat{e}_{i}^{2},$
@@
$\text{AIC}=2k+n\log(2\pi)+n\log(\sigma^{2})+\frac{1}{\sigma^{2}}\sum_{i=1}^{n}\hat{e}_{i}^{2}.$
@@ F.2 Test results
Appendix G Experiment Settings
@@
G.1 Judges, Candidates, and Tasks
@@
The reference (or baseline) answer-generating models are vicuna-13b-v1.3 for MTBench and human for DevBench. They are chosen to ensure a baseline quality of responses and an expected widely spread quality gap across evaluations. The other models that are compared to the reference models, namely “Model" in our context, are listed as follows.
@@
The model names are exactly what MTBench Zheng et al. (2024b) and DevBench Li et al. (2024a) used in their studies. That is why for GPTs, DevBench specifies the exact version (e.g., gpt-4-0613) while MTBench doesn’t (e.g., gpt-4). In this study, we directly use the provided answers of these models to the task questions to form answer pairs and queries for the LLM judges.
@@ G.2 Prompt Settings
  • Emphasized “should" and“shouldn’t"s.